# Cost-Effective Text Clustering with Large Language Models

### v1

@@ 5.2. Greedy Edge Selection
Proof.
Let $(t_{i},t_{j})=\underset{(t_{a},t_{b})\in\mathcal{S},\ a<b}{\operatorname{arg}\,\operatorname{max}}\;{a}$. In the extreme case, $j=i+1$ and only edges $(t_{a},t_{a+1}),(t_{a},t_{a+2}),\ldots,(t_{a},t_{j})$ are added into $\mathcal{S}$ for each $a\in[1,i-1]$ as they are surely larger than $(t_{i},t_{j})$. Put in another way, for each $t_{a}$ with $a\in[1,i]$, we can select $t_{b}$ with $a+1\leq b\leq j$, i.e., which corresponds to $j-a=i+1-a$ text instances. Thus, the total number of edges satisfy: $1+2+\ldots+i=\frac{(i+1)i}{2}=N$. Then, we can derive
$i^{2}+i-2N=0\Rightarrow i=\frac{\sqrt{8N+1}-1}{2},$
which finishes the proof. ∎
Specifically, Algorithm 1 begins by reordering the text instances in $\mathcal{T}$ according to their degrees defined in Eq. (2) in the ascending order, such that $\frac{1}{d(i)}$ of the $i$-th text $t_{i}$ is the $i$-th largest among all texts (Lines 1-2). Next, to facilitate the identification of the text $t_{b}$ that can yield the highest $\Phi(t_{a},t_{b})$ for each text $t_{a}$ ($b>a$), we maintain the index $p(t_{a})$ of the next suggested text for each possible $t_{a}$, which is initialized to be $a+1$ (Line 3). According to Lemma 5.2, the value of $a$ falls into the range of $[1,\lceil\frac{\sqrt{8N+1}-1}{2}\rceil]$. At Line 4, a stack $\mathcal{S}$ is additionally created to accommodate all the selected edges and $(t_{1},t_{2})$ is added to $\mathcal{S}$ as it has the globally largest SEC in $\mathcal{A}$. The index $p(t_{1})$ is then increased to $3$ since $t_{3}$ is the text that constitutes the new best edge with $t_{1}$ (Line 5). At a high level, the selection is to navigate between columns for the best edges in unvisited cells (in orange) as illustrated in Fig. 2(a).
Subsequently, Algorithm 1 starts an iterative process to greedily select the text pair $(t_{a},t_{b})$ ($b>a$) with the highest SEC to expand $\mathcal{S}$ (Lines 6-12). First, we get the latest added edge $(t_{a},t_{b})$ by picking the top element in stack $\mathcal{S}$ at Line 7. As such, the ideal edge $(t_{x},t_{y})$ to be selected in this round will be the one in $\mathcal{A}\setminus\mathcal{S}$ whose SEC is only second only to $\Phi(t_{a},t_{b})$. Simply, there are three possible choices for $(t_{x},t_{y})$, which satisfy $x=a$, $x>a$, or $x<a$, respectively. As per our greedy strategy, the first and second candidates are $(t_{a},t_{b+1})$ and $(t_{a+1},t_{b^{+}})$, where the index $b^{+}$ can be obtained via $p(t_{a+1})$. Intuitively, they are the two edges with the smallest SEC among all edges starting with $x=a$ and $x>a$, respectively. As for the case with $x<a$, Algorithm 1 greedily moves to the smallest value $x=1$ and selects $(t_{1},t_{b_{1}})$ ($b_{1}=p(t_{1})$) as the rest candidate (Line 8). We then form the set =$\mathcal{B}=\{(t_{1},t_{{b}^{-}}),(t_{a},t_{b+1}),(t_{a+1},t_{{b}^{+}})\}$ and find the one with the largest SEC as $(t_{x},t_{y})$, as in Lines 9-10. Next, at Lines 11-12, we push $(t_{x},t_{y})$ to $\mathcal{S}$ and update the index $p(t_{x})$ to $y+1$, i.e., the next best. Algorithm 1 repeats the above procedure until the size of $\mathcal{S}$ reaches $N$ (Line 6), and returns $\mathcal{S}$ as the final result.
@@
Edge Query Cluster CLINC docs by whether they are the same domain type. For each pair, respond with Yes or No without explanation.
@@
Given this context, do domain #1 and domain #2 likely correspond to the same general domain type?
Triangle Query Cluster CLINC docs by whether they are the same domain type. For each triangle, respond with a, b, c, d, or e without explanation.
@@
Given this context, do domain #1, domain #2 and domain #3 likely correspond to the same general domain type?
(a) All are same domain.  (b) Only #1 and #2 are same domain.
(c) Only #1 and #3 are same domain.  (d) Only #2 and #3 are same domain.  (e) None.
@@ 2 Push elements in $\mathcal{E}$ to $\mathcal{H}$;
6             if .$\mathcal{H}.\textsf{min}()\neq\beta+1$ then  Push $\beta$ to $\mathcal{H}$ and break ;
@@ 6.1. Spanning Triangle Centrality Maximization
Inspired by the SEC in Section 5.1, we define the spanning triangle centrality (STC) of any triangle $(t_{a},t_{b},t_{c})$ in $\mathcal{G}$ as
@@
Recall that $\Phi(t_{a},t_{b})$ signifies the fraction of spanning trees in $\mathcal{G}$ that include edge $(t_{a},t_{b})$. Analogously, $\Phi_{\triangle}(t_{a},t_{b},t_{c})$ can be roughly interpreted as the fraction of spanning trees of $\mathcal{G}$ where the triangle $(t_{a},t_{b},t_{c})$ appears. Using the of trick of estimating $\Phi(\cdot,\cdot)$ in Section 5.1, we reformulate $\Phi_{\triangle}(t_{a},t_{b},t_{c})$ as follows:
@@
Enumerating all possible triangles in $\mathcal{A}_{\triangle}$ for the optimization of Eq. (11) yields a computation cost of $O(n^{3})$, which is prohibitively expensive. Instead, akin to EdgeLLM, we greedily select the triangle $(t_{a},t_{b},t_{c})$ ($a<b<c$) with the largest $\Phi_{\triangle}(t_{a},t_{b},t_{c})$. To facilitate this, for each text $t_{a}\in\mathcal{T}$, we need to maintain a set $\mathcal{H}_{a}$ comprising text instances that occur with $t_{a}$ in triangles already selected. Then, to find the new triangle $(t_{a},t_{b},t_{c})$, we can simply fetch the smallest element from $[a+1,n]\setminus\mathcal{H}_{a}$ as $t_{b}$ and then from $[b+1,n]\setminus(\mathcal{H}_{a}\cup\mathcal{H}_{b})$ as $t_{c}$, respectively. Unfortunately, this approach requires materializing $\mathcal{H}_{a}$ consisting of many elements and sorting $[a+1,n]\setminus\mathcal{H}_{a}$, which is both space-consuming and computationally inefficient.
@@ 6.2. Greedy Triangle Selection
Complexity Analysis. As per our analysis in Section 5.2, Lines 1-2 consume $O(n\log{n}+nd)$ time. The time cost of each iteration at Lines 7-15 is dependent on the overhead for invoking GreedyScan and HeapAdd functions. Assume that the maximum size of the heap $\mathcal{H}_{i}\ \forall{i\in[1,n]}$ is $h$, then each call to GreedyScan is bounded by $O(h)$ as it scans all elements in the heap in the worst case. In HeapAdd, each insertion to $\mathcal{H}$ takes $O(\log{h})$ time, and each deletion consumes $O(1)$ time. In the worst case, HeapAdd will pop out $h-1$ elements. Then, each invocation of HeapAdd requires up to $O(h+\log{h})$ time. Considering $N_{\Delta}$ iterations, the overall time complexity is therefore $O(n\log{n}+nd+N_{\Delta}h)$. Due to Lines 3-6, the empirical value of $h$ is rather small and thus can be regarded as a constant.
@@ 6.3. Triangle Queries to LLMs
Subsequently, TriangleLLM requests for relations of each text triplet $(t_{a},t_{b},t_{c})$ in $\mathcal{S}_{\triangle}$ returned by Algorithm 4 from the LLM using the following prompting approach. Akin to the edge queries in Section 5.3, the input prompts specify the task context and contents of text instances $t_{a},t_{b},t_{c}$, whereas the query is framed as a multiple-choice question (MCQ). In particular, the MCQ includes a set of five valid alternatives as depicted in the l.h.s. of Figure 3, wherein a line between two text instances indicates a must-link while the empty space represents a cannot-link. Note that the r.h.s. three results are invalid as they lead to contradictions. For instance, when $(t_{a},t_{b})$ and $(t_{a},t_{c})$ are must-links, it can be inferred that $(t_{b},t_{c})$ should be a must-link, which contradicts the fact that $(t_{b},t_{c})$ is a cannot-link. We instruct the LLM to respond with one of the five valid options (a)-(e) in Table 1 and add the corresponding must-links and cannot-links to $\mathcal{M}$ and $\mathcal{C}$ accordingly.
@@ 7. Experiments
This section experimentally evaluates the effectiveness of the proposed TECL framework and its constituent techniques (EdgeLLM, TriangleLLM, WCSC, and WCKMeans) in enhancing text clustering quality. All experiments are conducted on a Linux machine with an NVIDIA A100 GPU (80GB RAM), AMD EPYC 7513 CPU (2.6 GHz), and 1TB RAM.
@@ 7.1. Experimental Seup
Datasets. Table  2 presents the statistics of a wide range of benchmarks for various tasks that are widely used for experimental evaluation in previous works (Zhang et al., 2023a; Viswanathan et al., 2023; Zhang et al., 2021a). The corpus size $\Omega$ represents the total number of text tokens of the text instances in the respective five text corpora, which is also the token budget we want to control in the main experiments in Table 3 for fair comparison. The BBC News  (Keraghel et al., 2024) dataset contains 2,225 headlines, categorized into 5 topics, while the Tweet (Yin and Wang, 2016) dataset consists of 2,472 tweets labeled for relevance to 89 queries from the Text Retrieval Conference [^2] https://trec.nist.gov/data/microblog.html. Bank77  (Zhang et al., 2021b) includes 3,080 customer service utterances classified into 77 intent categories. The Reddit  (Muennighoff et al., 2023) dataset, with 3,217 posts, is used for topic discovery, and CLINC (Zhang et al., 2022) contains 4,500 requests labeled by 10 domains.
@@
Baselines. We compare our proposed TECL against 10 baselines in text clustering. First, we include 7 embedding-based approaches that adopt 7 text embedders followed by K-Means++ or spectral clustering (Von Luxburg, 2007) for clustering, which include the classic TF-IDF, 4 popular PLMs: E5 (Wang et al., 2022), DistilBERT (Reimers, 2019), Sentence-BERT (Reimers, 2019), Instructor-Large (Su et al., 2023), and 2 LLMs: OpenAI-GPT (Brown et al., 2020) and LLaMA-2 (7B) (Touvron et al., 2023). Moreover, we consider three recent strong methods for text clustering: SCCL (Zhang et al., 2021a) utilizes contrastive learning to enhance cluster separation; ClusterLLM (Zhang et al., 2023a) leverages the feedback from an instruction-tuned LLM to fine-tune smaller embedding models and calibrate clustering granularity; PO-PCKMeans (Viswanathan et al., 2023) obtains pairwise constraints from LLMs with a subset of ground truth examples for semi-supervised clustering. For ClusterLLM, we also adopt its variant without fine-tuning for evaluation. The four versions of our TECL, i.e., EdgeLLM + WCKMeans, EdgeLLM + WCSC, TriangleLLM + WCKMeans, and TriangleLLM + WCSC, are included.
@@
By default, we utilize Instructor-Large as the primary text embedder in SCCL, ClusterLLM, PO-PCKMeans, and TECL, due to its consistently superior performance. Additionally, we employ GPT-4o-Mini as the backbone LLM. For a fair comparison, we set the budget $Q$ for querying LLMs to the corpus size $\Omega$ on each dataset in TECL, ClusterLLM, PO-PCKMeans, and the embedding-based methods OpenAI-GPT and LLaMA-2 (7B). The parameters of all competitors are set as suggested in their respective papers. Following prior studies (Zhang et al., 2023a; Viswanathan et al., 2023; Zhang et al., 2021a; Feng et al., 2024), clustering accuracy (ACC) and normalized mutual information (NMI) are used as the clustering quality metrics.
@@
Table 3. Clustering performance under the same query budget $Q=\Omega$ (best is highlighted in blue and runner-up underlined).
Method BBC News Tweet Bank77 Reddit CLINC
Embedding Clustering ACC NMI ACC NMI ACC NMI ACC NMI ACC NMI
TF-IDF K-means++ 24.88 1.84 55.19 77.62 36.35 58.83 12.87 18.66 25.38 14.09
SpectralClust 23.90 1.38 48.22 67.26 33.19 53.58 10.70 16.73 28.64 16.45
E5 K-means++ 71.66 53.03 60.90 85.91 58.74 77.09 38.88 48.75 59.60 56.25
SpectralClust 63.27 43.87 59.33 84.75 57.18 76.20 45.74 53.06 53.49 48.87
DistilBERT K-means++ 74.74 49.03 48.35 76.47 31.12 53.38 26.14 35.24 41.50 34.23
SpectralClust 71.42 41.82 51.08 77.62 39.71 60.45 30.40 38.44 35.47 27.79
Sentence-BERT K-means++ 79.02 56.88 62.74 86.42 60.77 78.84 42.31 50.02 50.32 48.98
SpectralClust 59.55 42.49 58.30 83.84 58.52 76.58 43.29 50.18 44.24 45.74
Instructor-Large K-means++ 86.53 71.51 65.34 87.74 64.57 82.01 53.62 62.62 55.94 57.41
SpectralClust 84.45 62.43 63.14 86.85 60.76 79.03 56.64 62.89 52.64 55.49
OpenAI-GPT K-means++ 88.56 73.65 65.42 87.64 64.20 81.74 51.06 60.10 55.31 56.05
SpectralClust 67.01 51.38 63.00 86.20 58.90 77.95 54.23 61.10 58.04 53.28
LLaMA-2 (7B) K-means++ 59.71 30.60 40.93 69.58 17.86 38.92 17.82 27.46 25.01 13.87
SpectralClust 49.05 24.11 52.29 76.27 27.60 49.33 24.43 34.67 27.08 18.46
SCCL (Zhang et al., 2021a) 83.60 62.00 36.50 68.50 36.00 61.30 28.70 36.00 46.60 44.00
ClusterLLM (Zhang et al., 2023a) 88.40 69.73 65.81 88.61 68.27 82.45 54.80 62.47 51.90 55.04
ClusterLLM w/o finetune (Zhang et al., 2023a) 89.07 70.97 64.59 87.78 65.14 82.33 54.75 62.87 58.66 58.95
PO-PCKMeans (Viswanathan et al., 2023) 89.53 72.16 65.58 88.20 67.79 81.92 55.10 62.51 57.80 56.32
TECL (EdgeLLM + WCKMeans) 90.47 73.96 71.16 88.90 65.29 82.52 57.07 63.80 62.31 61.11
TECL (EdgeLLM + WCSC) 86.88 69.49 70.91 88.51 67.99 82.55 57.01 63.21 56.20 57.77
TECL (TriangleLLM + WCKMeans) 90.65 74.25 72.25 88.60 66.72 83.07 59.28 64.98 65.16 62.09
TECL (TriangleLLM + WCSC) 91.01 75.22 73.10 89.08 69.03 83.20 57.78 64.25 57.91 60.02
Table 4. Ablation study on TECL.
Variant BBC News Tweet
ACC NMI ACC NMI
w/o WCSC 90.43 73.9 65.33 87.84
w/o TriangleLLM 88.72 70.25 67.27 87.44
TECL (TriangleLLM + WCSC) 91.01 75.22 73.10 89.08
TECLClusterLLMPO-PCKmeans
@@
${\Omega}/{10}$${\Omega}/{5}$${\Omega}/{2}$$\Omega$$2\Omega$0.620.640.660.680.700.720.74query budgetACC

### v2

@@ 5.2. Greedy Edge Selection
Specifically, Algorithm 1 begins by reordering the text instances in $\mathcal{T}$ according to their degrees defined in Eq. (2) in the ascending order, such that $\frac{1}{d(i)}$ of the $i$-th text $t_{i}$ is the $i$-th largest among all texts (Lines 1-2). Next, to facilitate the identification of the text $t_{b}$ that can yield the highest $\Phi(t_{a},t_{b})$ for each text $t_{a}$ ($b>a$), we maintain the index $p(t_{a})$ of the next suggested text for each possible $t_{a}$, which is initialized to be $a+1$ (Line 3). According to Lemma 5.2222The proof is deferred to Appendix A., the value of $a$ falls into the range of $1,\lceil\frac{\sqrt{8N+1}-1}{2}\rceil]$. At Line 4, a stack $\mathcal{S}$ is additionally created to accommodate all the selected edges and $(t_{1},t_{2})$ is added to $\mathcal{S}$ as it has the globally largest SEC in $\mathcal{A}$. The index $p(t_{1})$ is then increased to $3$ since $t_{3}$ is the text that constitutes the new best edge with $t_{1}$ (Line 5). At a high level, the selection is to navigate between columns for the best edges in unvisited cells (in orange) as illustrated in Fig. 2(a).
Subsequently, Algorithm 1 starts an iterative process to greedily select the text pair $(t_{a},t_{b})$ ($b>a$) with the highest SEC to expand $\mathcal{S}$ (Lines 6-12). First, we get the latest added edge $(t_{a},t_{b})$ by picking the top element in stack $\mathcal{S}$ at Line 7. As such, the ideal edge $(t_{x},t_{y})$ to be selected in this round will be the one in $\mathcal{A}\setminus\mathcal{S}$ whose SEC is only second only to $\Phi(t_{a},t_{b})$. Simply, there are three possible choices for $(t_{x},t_{y})$, which satisfy $x=a$, $x>a$, or $x<a$, respectively. As per our greedy strategy, the first and second candidates are $(t_{a},t_{b+1})$ and $(t_{a+1},t_{b^{+}})$, where the index $b^{+}$ can be obtained via $p(t_{a+1})$. Intuitively, they are the two edges with the smallest SEC among all edges starting with $x=a$ and $x>a$, respectively. As for the case with $x<a$, Algorithm 1 greedily moves to the smallest value $x=1$ and selects $(t_{1},t_{b_{1}})$ ($b_{1}=p(t_{1})$) as the rest candidate (Line 8). We then form the set $\mathcal{B}=\{(t_{1},t_{{b}^{-}}),(t_{a},t_{b+1}),(t_{a+1},t_{{b}^{+}})\}$ and find the one with the largest SEC as $(t_{x},t_{y})$, as in Lines 9-10. Next, at Lines 11-12, we push $(t_{x},t_{y})$ to $\mathcal{S}$ and update the index $p(t_{x})$ to $y+1$, i.e., the next best. Algorithm 1 repeats the above procedure until the size of $\mathcal{S}$ reaches $N$ (Line 6), and returns $\mathcal{S}$ as the final result.
@@
Edge Query Cluster CLINC docs by whether they are the same domain category. For each pair, respond with Yes or No without explanation.
@@
Given this context, do Domain #1 and Domain #2 likely correspond to the same domain category?
Triangle Query Cluster CLINC docs by whether they are the same domain category. For each triangle, respond with a, b, c, d, or e without explanation.
@@
Given this context, do Domain #1, Domain #2 and Domain #3 likely correspond to the same domain category?
(a) All are same category.  (b) Only #1 and #2 are same category.
(c) Only #1 and #3 are same category.  (d) Only #2 and #3 are same category.  (e) None.
@@ 2 Push elements in $\mathcal{E}$ to $\mathcal{H}$;
6     if $\mathcal{H}.\textsf{min}()\neq\beta+1$ then  Push $\beta$ to $\mathcal{H}$ and break ;
@@ 6.1. Spanning Triangle Centrality Maximization
Inspired by the SEC in § 5.1, we define the spanning triangle centrality (STC) of any triangle $(t_{a},t_{b},t_{c})$ in $\mathcal{G}$ as
@@
Recall that $\Phi(t_{a},t_{b})$ signifies the fraction of spanning trees in $\mathcal{G}$ that include edge $(t_{a},t_{b})$. Analogously, $\Phi_{\triangle}(t_{a},t_{b},t_{c})$ can be roughly interpreted as the fraction of spanning trees of $\mathcal{G}$ where the triangle $(t_{a},t_{b},t_{c})$ appears. Using the of trick of estimating $\Phi(\cdot,\cdot)$ in § 5.1, we reformulate $\Phi_{\triangle}(t_{a},t_{b},t_{c})$ as follows:
@@
Enumerating all possible triangles in $\mathcal{A}_{\triangle}$ for the optimization of Eq. (11) yields a computation cost of $O(n^{3})$, which is prohibitively expensive. Instead, akin to EdgeLLM, we greedily select the triangle $(t_{a},t_{b},t_{c})$ ($a<b<c$) with the largest $\Phi_{\triangle}(t_{a},t_{b},t_{c})$. To facilitate this, for each text $t_{a}\in\mathcal{T}$, we need to maintain a set $\mathcal{H}_{a}$ comprising text instances that occur with $t_{a}$ in triangles already selected. Then, to find the new triangle $(t_{a},t_{b},t_{c})$, we can simply fetch the smallest element from $a+1,n]\setminus\mathcal{H}_{a}$ as $t_{b}$ and then from $b+1,n]\setminus(\mathcal{H}_{a}\cup\mathcal{H}_{b})$ as $t_{c}$, respectively. Unfortunately, this approach requires materializing $\mathcal{H}_{a}$ consisting of many elements and sorting $a+1,n]\setminus\mathcal{H}_{a}$, which is both space-consuming and computationally inefficient.
@@ 6.2. Greedy Triangle Selection
Complexity Analysis. As per our analysis in § 5.2, Lines 1-2 consume $O(n\log{n}+nd)$ time. The time cost of each iteration at Lines 7-15 is dependent on the overhead for invoking GreedyScan and HeapAdd functions. Assume that the maximum size of the heap $\mathcal{H}_{i}\ \forall{i\in[1,n]}$ is $h$, then each call to GreedyScan is bounded by $O(h)$ as it scans all elements in the heap in the worst case. In HeapAdd, each insertion to $\mathcal{H}$ takes $O(\log{h})$ time, and each deletion consumes $O(1)$ time. In the worst case, HeapAdd will pop out h1h-1 elements. Then, each invocation of HeapAdd requires up to $O(h+\log{h})$ time. Considering $N_{\Delta}$ iterations, the overall time complexity is therefore $O(n\log{n}+nd+N_{\Delta}h)$. Due to Lines 3-6, the empirical value of $h$ is rather small and thus can be regarded as a constant.
@@ 6.3. Triangle Queries to LLMs
Subsequently, TriangleLLM requests for relations of each text triplet $(t_{a},t_{b},t_{c})$ in $\mathcal{S}_{\triangle}$ returned by Algorithm 4 from the LLM using the following prompting approach. Akin to the edge queries in § 5.3, the input prompts specify the task context and contents of text instances $t_{a},t_{b},t_{c}$, whereas the query is framed as a multiple-choice question (MCQ). In particular, the MCQ includes a set of five valid alternatives as depicted in the l.h.s. of Fig. 3, wherein a line between two text instances indicates a must-link while the empty space represents a cannot-link. Note that the r.h.s. three results are invalid as they lead to contradictions. For instance, when $(t_{a},t_{b})$ and $(t_{a},t_{c})$ are must-links, it can be inferred that $(t_{b},t_{c})$ should be a must-link, which contradicts the fact that $(t_{b},t_{c})$ is a cannot-link. We instruct the LLM to respond with one of the five valid options (a)-(e) in Table 1 and add the corresponding must-links and cannot-links to $\mathcal{M}$ and $\mathcal{C}$ accordingly.
@@ 7. Experiments
This section experimentally evaluates the effectiveness of the proposed Cequel framework and its constituent techniques (EdgeLLM, TriangleLLM, WCSC, and WCKMeans) in enhancing text clustering quality. All experiments are conducted on a Linux machine with an NVIDIA A100 GPU (80GB RAM), AMD EPYC 7513 CPU (2.6 GHz), and 1TB RAM. The source code and datasets are publicly accessible at https://github.com/HKBU-LAGAS/Cequel.
@@ 7.1. Experimental Seup
Datasets. Table  2 presents the statistics of a wide range of benchmarks for various tasks that are widely used for experimental evaluation in previous works (Zhang et al., 2023a; Viswanathan et al., 2023; Zhang et al., 2021a). The corpus size $\Omega$ represents the total number of text tokens of the text instances in the respective five text corpora, which is also the token budget we want to control in the main experiments in Table 3 for fair comparison. We evaluate six benchmark datasets: BBC News (2,225 texts, 5 topics) (Keraghel et al., 2024), Tweet (2,472 tweets, 89 TREC [^3] https://trec.nist.gov/data/microblog.html queries) (Yin and Wang, 2016), Bank77 (3,080 utterances, 77 intents) (Zhang et al., 2021b), Reddit (3,217 posts, 50 topics) (Muennighoff et al., 2023), CLINC (4,500 requests, 10 domains) (Zhang et al., 2022), and Massive (11,514 utterances, 18 scenarios) (FitzGerald et al., 2022).
@@
Baselines. We compare our proposed Cequel against 10 baselines in text clustering. First, we include 7 embedding-based approaches that adopt 7 text embedders followed by K-Means++ or spectral clustering (Von Luxburg, 2007) for clustering, which include the classic TF-IDF, 4 popular PLMs: E5 (Wang et al., 2022), DistilBERT (Reimers, 2019), Sentence-BERT (Reimers, 2019), Instructor-Large (Su et al., 2023), and 2 LLMs: OpenAI-GPT (Brown et al., 2020) and LLaMA-2 (7B) (Touvron et al., 2023). Moreover, we consider three recent strong methods for text clustering: SCCL (Zhang et al., 2021a) utilizes contrastive learning to enhance cluster separation; ClusterLLM (Zhang et al., 2023a) leverages the feedback from an instruction-tuned LLM to fine-tune smaller embedding models and calibrate clustering granularity; PO-PCKMeans (Viswanathan et al., 2023) obtains pairwise constraints from LLMs with a subset of ground truth examples for semi-supervised clustering. For ClusterLLM, we also adopt its variant without fine-tuning for evaluation. The four versions of our Cequel, i.e., EdgeLLM + WCKMeans, EdgeLLM + WCSC, TriangleLLM + WCKMeans, and TriangleLLM + WCSC, are included.
@@
By default, we utilize Instructor-Large as the primary text embedder in SCCL, ClusterLLM, PO-PCKMeans, and Cequel, due to its consistently superior performance. Additionally, we employ GPT-4o-Mini as the backbone LLM. For a fair comparison, we set the budget $Q$ for querying LLMs to the corpus size $\Omega$ on each dataset in Cequel, ClusterLLM, PO-PCKMeans, and the embedding-based method OpenAI-GPT. The parameters of all competitors are set as suggested in their respective papers. Following prior studies (Zhang et al., 2023a; Viswanathan et al., 2023; Zhang et al., 2021a; Feng et al., 2024), clustering accuracy (ACC) and normalized mutual information (NMI) are used as the clustering quality metrics.
@@
Table 3. Clustering performance under the same query budget $Q=\Omega$. Best results are bolded, while best baselines are underlined.
Method BBC News Tweet Bank77 Reddit CLINC Massive
Embedding Clustering ACC NMI ACC NMI ACC NMI ACC NMI ACC NMI ACC NMI
TF-IDF K-means++ 24.88 1.84 55.19 77.62 36.35 58.83 12.87 18.66 25.38 14.09 32.03 30.00
SpectralClust 23.90 1.38 48.22 67.26 33.19 53.58 10.70 16.73 28.64 16.45 34.15 29.35
E5 K-means++ 71.66 53.03 60.90 85.91 58.74 77.09 38.88 48.75 59.60 56.25 41.56 42.36
SpectralClust 63.27 43.87 59.33 84.75 57.18 76.20 45.74 53.06 53.49 48.87 46.05 46.76
DistilBERT K-means++ 74.74 49.03 48.35 76.47 31.12 53.38 26.14 35.24 41.50 34.23 30.04 28.13
SpectralClust 71.42 41.82 51.08 77.62 39.71 60.45 30.40 38.44 35.47 27.79 29.23 29.51
Sentence-BERT K-means++ 79.02 56.88 62.74 86.42 60.77 78.84 42.31 50.02 50.32 48.98 59.21 64.17
SpectralClust 59.55 42.49 58.30 83.84 58.52 76.58 43.29 50.18 44.24 45.74 54.74 60.69
Instructor-Large K-means++ 86.53 71.51 65.34 87.74 64.57 82.01 53.62 62.62 55.94 57.41 59.19 66.21
SpectralClust 84.45 62.43 63.14 86.85 60.76 79.03 56.64 62.89 52.64 55.49 58.53 60.63
OpenAI-GPT K-means++ 88.56 73.65 65.42 87.64 64.20 81.74 51.06 60.10 55.31 56.05 62.28 68.08
SpectralClust 67.01 51.38 63.00 86.20 58.90 77.95 54.23 61.10 58.04 53.28 60.76 63.16
LLaMA-2 (7B) K-means++ 59.71 30.60 40.93 69.58 17.86 38.92 17.82 27.46 25.01 13.87 18.92 15.17
SpectralClust 49.05 24.11 52.29 76.27 27.60 49.33 24.43 34.67 27.08 18.46 25.61 21.20
SCCL (Zhang et al., 2021a) 83.60 62.00 36.50 68.50 36.00 61.30 28.70 36.00 46.60 44.00 45.90 45.93
ClusterLLM (Zhang et al., 2023a) 88.40 69.73 65.81 88.61 68.27 82.45 54.80 62.47 51.90 55.04 61.02 67.92
ClusterLLM w/o finetune (Zhang et al., 2023a) 89.07 70.97 64.59 87.78 65.14 82.33 54.75 62.87 58.66 58.95 59.37 66.60
PO-PCKMeans (Viswanathan et al., 2023) 89.53 72.16 65.58 88.20 67.79 81.92 55.10 62.51 57.80 56.32 64.11 65.99
Cequel (EdgeLLM + WCKMeans) 90.47 73.96 71.16 88.90 65.29 82.52 57.07 63.80 62.31 61.11 67.66 68.16
Cequel (EdgeLLM + WCSC) 86.88 69.49 70.91 88.51 67.99 82.55 57.01 64.25 56.20 57.77 60.50 66.05
Cequel (TriangleLLM + WCKMeans) 90.65 74.25 72.25 88.60 66.72 83.07 59.28 64.98 65.16 62.09 69.13 70.44
Cequel (TriangleLLM + WCSC) 91.01 75.22 73.10 89.08 69.03 83.20 57.78 64.25 57.91 60.02 61.51 67.05
Table 4. Ablation study on Cequel.
Variant BBC News Tweet Bank77 Reddit
ACC NMI ACC NMI ACC NMI ACC NMI
Cequel 90.47 73.96 71.16 88.90 65.29 82.52 57.07 63.80
w/o WCKMeans 88.65 71.82 69.30 87.95 63.25 81.25 55.48 62.15
w/o EdgeLLM 89.12 72.35 68.85 87.67 64.13 81.85 56.12 62.97
Cequel 86.88 69.49 70.91 88.51 67.99 82.55 57.01 64.25
w/o WCSC 85.21 67.94 68.25 87.68 66.44 81.71 55.52 63.24
w/o EdgeLLM 85.80 68.55 67.91 87.48 66.85 81.95 56.32 63.45
Cequel 90.65 74.25 72.25 88.60 66.72 83.07 59.28 64.98
w/o WCKMeans 89.80 72.38 65.66 87.84 64.85 82.17 57.22 64.12
w/o TriangleLLM 88.25 71.60 69.53 87.81 65.16 81.75 56.57 63.59
Cequel 91.01 75.22 73.10 89.08 69.03 83.20 57.78 64.25
w/o WCSC 90.43 73.90 65.33 87.84 67.12 82.65 56.30 63.74
w/o TriangleLLM 88.72 70.25 67.27 87.44 66.49 83.10 57.22 63.83
CequelClusterLLMPO-PCKmeans
@@
${\Omega}/{10}$${\Omega}/{5}$${\Omega}/{2}$$\Omega$$2\Omega$0.590.620.650.680.710.74query budgetACC
@@
${\Omega}/{10}$${\Omega}/{5}$${\Omega}/{2}$$\Omega$$2\Omega$0.650.660.670.680.690.70query budgetACC
((c)) Bank77
${\Omega}/{10}$${\Omega}/{5}$${\Omega}/{2}$$\Omega$$2\Omega$0.530.540.550.560.570.58query budgetACC
((d)) Reddit